Write an equation in Standard Form such that it represents with slope -8/3 and it passes (1/2,-3)
Answer by lwsshak3(11628)

Write an equation in Standard Form such that it represents with slope -8/3 and it passes (1/2,-3)
***
Standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
Equation: y=-8x/3+b
solve for b using given coordinates(1/2,-3) on the line
-3=-(8*1/2)/3+b
-3=-4/3+b
b=-5/3
equation: y=-8x/3-5/3